You bought 12 2-liter bottles of regular and diet pop for a school dance. The regular pop was on sale for $1.00 per 2-liter bottle. The diet pop cost $1.50 per 2-liter bottle. You spent $15.00. How many of each type of pop did you buy?
Answer by stanbon(48516)

Let number of $1.00 bottles be "x"
Then number of $1.50 bottles must be "12-x"
Value of $1.00 bottles is (1)x
Value of $2.00 bottles is (2)(12-x)=24-2x
EQUATION:
Value + Value = 15 dollars
x+24-2x=15
-x=-9
x=9 (number of $1.00 bottles)
12-x=3 (number of $2.00 bottles)
Cheers,
Stan H.